the sum of a certain even no. and the fourth even no. after it is 48. find the no.
Answer by edjones(7311)

Let x, x+8 be the numbers
x+x+8=48
2x+8=48
subtract 8 from each side: 2x=40
divide 2 into each side: x=20
Check
x+8=28
20+28=48
Ed